# Investigating the Impact of Project Risks on Employee Turnover Intentions in the IT Industry of Pakistan

### v3

@@ 7 Measurement Of Variables
We measured all the items using 5 points Likert scale Wuensch (2005). The scale developed by Arpita Sharma, and Aauushi in their research paper measures Project-specific risk factors. Similarly, the three-question scale is defined to measure Job satisfaction  Cammann (1979) Price and Muller scale analyze Perceived job alternatives  Mobley (1977). External networking and social networking impact on perceived job alternatives is computed using the scale developed by  Scott M. Soltis (2013) and finally employee turnover intention is measured by a scale defined by Michaels (1987).
@@ 8 Data reliability
The data collected from IT professionals reflects their individual experiences in the industry, likely resulting in varied responses. To ensure the reliability of our scales, we conducted a Cronbach alpha score test LJ (1951). The Cronbach alpha score ranges from 0 to 1, with values exceeding 0.7 considered good. Table 5 presents the Cronbach alpha scores of our scales
@@ 9.1 Correlation Analysis
Pearson correlation measures the linear relation in two values ’x’ and y. The value of that correlation varies between -1 and 1. The higher positive values show the positive correlation between two variables, while values closer to -1 indicate a stronger negative correlation. In our experiments, we computed the correlation between project-specific risk factors and job satisfaction. Table 6 displays the results of the correlation analysis involving project-specific risk factors, perceived job alternatives, and job satisfaction.
Table 6: Correlational of project specific risk factors with job satisfaction
@@
In our analysis, we observed that team composition risk had a positive impact on job satisfaction, indicating that addressing issues related to team composition led to greater job satisfaction among employees. Similarly, addressing risks associated with software requirements estimation positively influenced job satisfaction, contributing to higher levels of satisfaction among employees. However, the impact of control process risk on job satisfaction was found to be non-significant, suggesting that it did not affect job satisfaction significantly. On the other hand, the effect of dependability risk on job satisfaction was significant and positive.
@@
Job satisfaction has a significant negative relationship with turnover intentions leading to fewer turnover intentions when people are happier with their work. Similarly, perceived job opportunities have a non-significant relation with turnover intentions. In the final set of experiments of Pearson’s correlation, we computed the relationship between internal and external networking and perceived job alternatives. Table 8 shows our results.
@@ 9.2 Direct Regression Analysis
Initially, control values were entered into the model, revealing that they did not significantly affect job satisfaction, with a p-value higher than 0.05. The R square value of control variables indicated they accounted for a 4.1% variance in job satisfaction. Next, we added the team composition risk factor to the model. In the presence of team composition risk, control variables became significantly related to job satisfaction, explaining an additional 47.2% variance in job satisfaction. Reductions in team composition risks positively impacted job satisfaction, with a coefficient (B) of 0.622 and p-value < 0.05. Overall, the analysis provided support for hypotheses 1(a), 1(c), and 1(d), indicating that team composition risks, dependability risks, and risks involved in estimating software requirements had significant effects on job satisfaction. However, control process risks did not significantly affect job satisfaction.
@@
Control process risk did not show significant correlation with job satisfaction (Beta = 0.46, p > 0.05), suggesting that control process risks are not significantly related to job satisfaction. In the case of dependability risk (fourth column of the table), none of the control variables had a significant impact on job satisfaction in the presence of dependability risk. However, reducing dependability risk positively influenced job satisfaction (Beta = 0.499, p < 0.05). Dependability risks explained an additional 35.9% variance in job satisfaction.
@@
When considering risks involved in estimating software requirements (fifth column of the table), control variables did not show significant correlation with job satisfaction. However, risks involved in estimating software requirements had a positive influence on job satisfaction (Beta = 0.578, p < 0.05). These risks explained 42.8% of the variance in job satisfaction.
@@
Thus, hypotheses 1(a), 1(c), and 1(d) are supported, indicating that team composition risks, dependability risks, and risks involved in estimating software requirements significantly impact job satisfaction. However, hypothesis 1(b) is not supported, suggesting that control process risks do not have a significant effect on job satisfaction.
@@
Furthermore, job satisfaction was found to have a significant negative relationship with turnover intention (Beta = -0.560, p < 0.01), explaining 42% of the variance in employee turnover intention. Moreover, perceived job alternatives were positively associated with turnover intentions, accounting for an additional 51% of the variance in turnover intentions. The beta coefficient value of perceived job alternatives was 0.69, indicating a significant positive relation with turnover intentions (p < 0.01).
@@
To evaluate hypotheses 4(a) and 4(b), we regressed external networking and internal networking against perceived job alternatives. Internal networking did not influence perceived job alternatives significantly (p > 0.05), with an R square value of only 1.4%, indicating that internal networking explains very little variance in perceived job alternatives. Similarly, external social networking did not have a significant influence on perceived job alternatives, with a low R square value suggesting minimal impact on employee turnover intentions.
@@ 9.3 Moderating Effect of perceived Job Alternatives:
In step 1 of our experiment, we introduced job satisfaction as the independent variable. Job satisfaction exhibited a significant negative effect on turnover intention (B = -0.565, p < 0.01), explaining 43% of the variance in employee turnover intention.
In step 2, we added perceived job alternatives to the model. The regression coefficient for perceived job alternatives was found to be B = 0.714, indicating a significant positive impact on turnover intention. Perceived job alternatives explained an additional 52% of the variance in turnover intention. Consequently, the predictive capacity of our model improved with the inclusion of perceived job alternatives. Together, job satisfaction and perceived job alternatives accounted for 58% of the variance in turnover intentions. In step 3, we introduced an interaction term into the model by multiplying job satisfaction and perceived job alternatives. The purpose was to assess whether perceived job alternatives moderate the relationship between job satisfaction and turnover intention. However, the results showed that the interaction term had a non-significant impact (B = 0.095, p > 0.05), indicating that perceived job alternatives do not moderate the relationship between job satisfaction and employee turnover. Therefore, hypothesis (2) of our study is not supported.
@@
R square -.565 0.65 0.387
** denotes significant two tail test where p< 0.01 * denotes the 1 tail test where p < 0.05
9.4 Meditation Test
To evaluate hypothesis (3), we analyzed whether job satisfaction fully or partially mediates the association between project-specific risk factors and turnover intention. In step 1 of our analysis, control process risk was found to have no significant direct association with turnover intention (beta = -0.21, p > 0.05). In step 2, control process risk was regressed on job satisfaction, revealing a non-significant relationship (beta = 0.19, p > 0.01). However, in step 3, job satisfaction was negatively associated with turnover intentions (beta = -0.561, p < 0.01).
@@ 9.4 Meditation Test
R square 0.017 0.022 0.48 0.114
Dependability risk has a significant impact on turnover intentions (Beta = -0.501, p < 0.05). In step 2, we examined the relationship between Dependability risk and job satisfaction, finding that Dependability risks positively influence job satisfaction (Beta = 0.498, p < 0.01). Step 3 revealed that job satisfaction significantly affects turnover intentions (beta = -0.582, p < 0.01). In step four of the experiment, job satisfaction was used as a control variable. Multiple regression analysis showed that both Dependability risks (beta = -0.452, p < 0.01) and job satisfaction (beta = -0.356, p < 0.01) impact turnover intention. This suggests that job satisfaction partially mediates the relationship between Dependability risk and turnover intentions.
@@
R square 0.369 0.378 0.49 0.46
RIESR also has a significant impact on turnover intentions, with a Beta value of 0.578 and p < 0.01. In step 2 of our experiment, we included job satisfaction in the model to assess the association between RIESR and job satisfaction. Similarly, RIESR is significantly related to job satisfaction (Beta = 0.547, p < 0.01). In step 3, we examined the relationship between job satisfaction and turnover intentions. Job satisfaction was found to be negatively related to turnover intentions (Beta = -0.596, p < 0.01). In step 4, job satisfaction was used as a control variable. Multiple regression analysis revealed that both RIESR (Beta = 0.571, p < 0.01) and job satisfaction (Beta = 0.456, p < 0.01) impact turnover. This indicates that satisfaction with a job partially mediates the association between RIESR and turnover intentions.
@@
R square 0.369 0.378 0.49 0.46
In the final experiment, we examined whether Team Risks mediate through job satisfaction and significantly influence employee turnover. In step 1, we assessed the influence of Team Risks on turnover intentions. Team risk was found to have a significant impact on turnover intentions (Beta = -0.512, p < 0.01). In step 2, we introduced job satisfaction into the model. Team risks were significantly related to job satisfaction (Beta = 0.642, p < 0.01). In step 3, we analyzed the influence of job satisfaction on turnover intentions (Beta = 0.578, p < 0.01). Finally, in step 4, job satisfaction was included as a control variable. Multiple regression analysis revealed that both team composition risk (Beta = 0.593, p < 0.01) and job satisfaction (Beta = 0.481, p < 0.01) significantly influence turnover intentions. This suggests that job satisfaction partially mediates the association between team composition risk and turnover intentions.
@@ 10 Conclusions
The research study delves into the impact of project-specific risks on turnover intentions, the role of external and internal networking in perceived job alternatives in Pakistan’s IT industry, and the development of a classification model to predict employee turnover intentions. This exploration is particularly novel in the context of the IT industry in Pakistan, which plays a crucial role in driving the country’s economy and creating employment opportunities. Given that IT professionals often juggle multiple projects simultaneously, they encounter various project-specific risks. Understanding these risks and their effects is vital for project success and mitigating turnover intentions. This study aimed to investigate the influence of project risks on turnover intentions, mediated by job satisfaction, as well as to explore the moderating role of perceived job alternatives (PJA). Additionally, the study sought to examine the impact of external and internal networking on PJA, alongside the overarching objective of developing a predictive model for employee turnover intentions.
To gather data, we distributed a questionnaire to numerous companies across Pakistan, resulting in responses from 150 participants. Our analyses encompassed various statistical techniques, including regression analysis, step-wise linear regression, and correlation analysis, to evaluate our hypotheses. Our findings revealed that minimizing project-specific risks is essential for reducing turnover intentions. Specifically, all project-specific risk factors, with the exception of control process risk, mediated through job satisfaction and influenced employee turnover. Additionally, while perceived job alternatives did not moderate the relationship between job satisfaction and turnover intention, they independently influenced turnover intention. Similarly, our analyses indicated that external and internal networking did not directly impact turnover intention. In summary, this study sheds light on critical factors influencing turnover intentions in Pakistan’s IT industry and underscores the importance of managing project-specific risks and perceived job alternatives to mitigate turnover.
@@ 11 Future Work
In the future, we plan to expand our research to explore project-specific risks across diverse industries, such as construction companies. By investigating the factors contributing to project-specific risks in various sectors, we aim to gain a comprehensive understanding of their influence on employee turnover intentions. Additionally, we intend to delve deeper into the determinants of employee turnover intentions, examining a wide range of influencing factors. This comprehensive approach will enable us to uncover nuanced insights into the drivers of turnover intentions and their implications for organizational management. Moreover, we recognize the importance of considering cultural differences in our research. We aim to conduct cross-cultural analyses to explore how cultural norms and values shape employee turnover intentions. By collecting data from different cultural contexts, we can better understand the variations in turnover intentions across diverse populations.
@@
Furthermore, we plan to evaluate the effectiveness of organizational initiatives and activities in mitigating employee turnover intentions. By assessing the impact of various organizational interventions, such as training programs, leadership initiatives, and employee engagement strategies, we can identify effective measures for reducing turnover and retaining valuable talent. Overall, our future research endeavors will contribute to the development of evidence-based strategies for minimizing employee turnover intentions and enhancing organizational effectiveness, particularly in dynamic industries like IT where talent retention is crucial.

### v4

@@ 7 Measurement Of Variables
We measured all the items using a 5-point Likert scale 75323:2931936. The scale proposed by Arpita Sharma, and Aauushi in their research paper measures Project-specific risk factors. Similarly, the three-question scale is defined to measure Job satisfaction  75323:2894280 Price and Muller scale analyze Perceived job alternatives  75323:2894203. External networking and social networking impact on perceived job alternatives is computed using the scale proposed by  75323:2894281 and finally employee turnover intention is measured by a scale defined by 75323:2894282.
@@ 8 Data reliability
The data collected from IT professionals reflects their individual experiences in the industry, likely resulting in varied responses. To aim to the reliability of our scales, we conducted a Cronbach alpha score test 75323:2931937. The Cronbach alpha score ranges from 0 to 1, with values exceeding 0.7 considered good. Table 5 presents the Cronbach alpha scores of our scales
@@ 9.1 Correlation Analysis
Pearson correlation measures the linear relationship between two variables x and y. The value of that correlation varies between -1 and 1. The higher positive values show the positive correlation between two variables, while values closer to -1 indicate a stronger negative correlation. In our experiments, we computed the correlation between project-specific risk factors and job satisfaction. Table 6 displays the results of the correlation analysis involving project-specific risk factors, perceived job alternatives, and job satisfaction.
Table 6: Correlation of project specific risk factors with job satisfaction
@@
In our analysis, we observed that team composition risk was positively correlated/associated with job satisfaction, indicating that addressing issues related to team composition was positively correlated with job satisfaction among employees. Similarly, addressing risks associated with software requirements estimation was positively correlated with job satisfaction, contributing to higher levels of satisfaction among employees. However, the association between control process risk and job satisfaction was not statistically significant, suggesting that it did not affect job satisfaction significantly. On the other hand, the effect of dependability risk on job satisfaction was notable and positive.
@@
Job satisfaction was associated with lower turnover intentions leading to fewer turnover intentions when people are happier with their work. Similarly, perceived job opportunities have no statistically significant relationship with turnover intentions. In the final set of experiments of Pearson’s correlation, we computed the relationship between internal and external networking and perceived job alternatives. Table 8 shows our results.
@@ 9.2 Direct Regression Analysis
Initially, control values were entered into the model, revealing that they did not significantly affect job satisfaction, with a p-value higher than 0.05. The $R^{2}$ value of control variables indicated they accounted for a 4.1% variance in job satisfaction. Next, we added the team composition risk factor to the model. In the presence of team composition risk, control variables became significantly related to job satisfaction, explaining an additional 47.2% variance in job satisfaction. Reductions in team composition risks was significantly positively associated with job satisfaction, with a coefficient (B) of 0.622 and p-value < 0.05. Overall, the analysis was consistent with the hypotheses 1(a), 1(c), and 1(d), indicating that team composition risks, dependability risks, and risks involved in estimating software requirements had notable effects on job satisfaction. However, control process risks did not significantly affect job satisfaction.
@@
Control process risk did not show notable correlation with job satisfaction (Beta = 0.46, p > 0.05), suggesting that control process risks are not significantly related to job satisfaction. In the case of dependability risk (fourth column of the table), none of the control variables showed a notable association with job satisfaction in the presence of dependability risk. However, reducing dependability risk was positively associated with job satisfaction (Beta = 0.499, p < 0.05). Dependability risks explained an additional 35.9% variance in job satisfaction.
@@
When considering risks involved in estimating software requirements (fifth column of the table), control variables did not show notable correlation with job satisfaction. However, risks involved in estimating software requirements had a positive influence on job satisfaction (Beta = 0.578, p < 0.05). These risks explained 42.8% of the variance in job satisfaction.
@@
Thus, hypotheses 1(a), 1(c), and 1(d) are supported, indicating that team composition risks, dependability risks, and risks involved in estimating software requirements significantly impact job satisfaction. However, hypothesis 1(b) is not supported, suggesting that control process risks do not have a notable effect on job satisfaction.
@@
Furthermore, job satisfaction was found to have a notable negative relationship with turnover intention (Beta = -0.560, p < 0.01), explaining 42% of the variance in employee turnover intention. Moreover, perceived job alternatives were positively associated with turnover intentions, accounting for an additional 51% of the variance in turnover intentions. The beta coefficient value of perceived job alternatives was 0.69 (p < 0.01) in the regression model (Table 10); however, Table 7 shows no significant bivariate correlation between perceived job alternatives and turnover intentions (r = 0.16), so this result should be interpreted as conditional on the regression model.
@@
To evaluate hypotheses 4(a) and 4(b), we regressed external networking and internal networking against perceived job alternatives. Internal networking did not influence perceived job alternatives significantly (p > 0.05), with an $R^{2}$ value of only 1.4%, indicating that internal networking explains very little variance in perceived job alternatives. Similarly, external social networking did not have a notable influence on perceived job alternatives, with a low $R^{2}$ value suggesting minimal impact on employee turnover intentions.
@@ 9.3 Moderating Effect of perceived Job Alternatives:
In step 1 of our experiment, we introduced job satisfaction as the independent variable. Job satisfaction exhibited a notable negative effect on turnover intention (B = -0.565, p < 0.01), explaining 43% of the variance in employee turnover intention.
In step 2, we added perceived job alternatives to the model. The regression coefficient for perceived job alternatives was found to be B = 0.714, indicating a notable positive impact on turnover intention. Perceived job alternatives explained an additional 52% of the variance in turnover intention. Consequently, the predictive capacity of our model improved with the inclusion of perceived job alternatives. Together, job satisfaction and perceived job alternatives accounted for 58% of the variance in turnover intentions. In step 3, we introduced an interaction term into the model by multiplying job satisfaction and perceived job alternatives. The purpose was to assess whether perceived job alternatives moderate the relationship between job satisfaction and turnover intention. However, the results showed that the interaction term had a not statistically significant impact (B = 0.095, p > 0.05), indicating that perceived job alternatives do not moderate the relationship between job satisfaction and employee turnover. Therefore, hypothesis (3) of our study is not supported.
@@
$R^{2}$ -.565 0.65 0.387
** denotes notable two tail test where p< 0.01 * denotes the 1 tail test where p < 0.05
9.4 Mediation Test
To evaluate hypothesis (3), we analyzed whether job satisfaction fully or partially mediates the association between project-specific risk factors and turnover intention. In step 1 of our analysis, control process risk was found to have no notable direct association with turnover intention (beta = -0.21, p > 0.05). In step 2, control process risk was regressed on job satisfaction, revealing a not statistically significant relationship (beta = 0.19, p > 0.01). However, in step 3, job satisfaction was negatively associated with turnover intentions (beta = -0.561, p < 0.01).
@@ 9.4 Mediation Test
$R^{2}$ 0.017 0.022 0.48 0.114
Dependability risk has a notable impact on turnover intentions (Beta = -0.501, p < 0.05). In step 2, we examined the relationship between Dependability risk and job satisfaction, finding that Dependability risks was positively correlated with job satisfaction (Beta = 0.498, p < 0.01). Step 3 revealed that job satisfaction significantly affects turnover intentions (beta = -0.582, p < 0.01). In step four of the experiment, job satisfaction was used as a control variable. Multiple regression analysis showed that both Dependability risks (beta = -0.452, p < 0.01) and job satisfaction (beta = -0.356, p < 0.01) impact turnover intention. This suggests that job satisfaction partially mediates the relationship between Dependability risk and turnover intentions.
@@
$R^{2}$ 0.369 0.378 0.49 0.46
Risks involved in estimating software requirements (RIESR) also has a notable impact on turnover intentions, with a Beta value of 0.578 and p < 0.01. In step 2 of our experiment, we included job satisfaction in the model to assess the association between RIESR and job satisfaction. Similarly, RIESR is significantly related to job satisfaction (Beta = 0.547, p < 0.01). In step 3, we examined the relationship between job satisfaction and turnover intentions. Job satisfaction was found to be negatively related to turnover intentions (Beta = -0.596, p < 0.01). In step 4, job satisfaction was used as a control variable. Multiple regression analysis revealed that both RIESR (Beta = 0.571, p < 0.01) and job satisfaction (Beta = 0.456, p < 0.01) impact turnover. This indicates that satisfaction with a job partially mediates the association between RIESR and turnover intentions.
@@
$R^{2}$ 0.369 0.378 0.49 0.46
In the final experiment, we examined whether Team Risks mediate through job satisfaction and significantly influence employee turnover. In step 1, we assessed the influence of Team Risks on turnover intentions. Team risk was found to have a notable impact on turnover intentions (Beta = -0.512, p < 0.01). In step 2, we introduced job satisfaction into the model. Team risks were significantly related to job satisfaction (Beta = 0.642, p < 0.01). In step 3, we analyzed the influence of job satisfaction on turnover intentions (Beta = 0.578, p < 0.01). Finally, in step 4, job satisfaction was included as a control variable. Multiple regression analysis revealed that both team composition risk (Beta = 0.593, p < 0.01) and job satisfaction (Beta = 0.481, p < 0.01) significantly influence turnover intentions. This suggests that job satisfaction partially mediates the association between team composition risk and turnover intentions.
@@ 10 Conclusions
The research study delves into the impact of project-specific risks on turnover intentions, the role of external and internal networking in perceived job alternatives in Pakistan’s IT industry, and the development of a statistical analysis to examine employee turnover intentions. This exploration is particularly context-specific in the context of the IT industry in Pakistan, which plays a relevant role in driving the country’s economy and creating employment opportunities. Given that IT professionals often juggle multiple projects simultaneously, they encounter various project-specific risks. Understanding these risks and their effects is vital for project success and mitigating turnover intentions. This study aimed to investigate the influence of project risks on turnover intentions, mediated by job satisfaction, as well as to explore the moderating role of perceived job alternatives (PJA). Additionally, the study sought to examine the impact of external and internal networking on PJA, alongside the main objective of developing an analytical framework for employee turnover intentions.
To gather data, we distributed a questionnaire to numerous companies across Pakistan, resulting in responses from 150 participants. Our analyses encompassed various statistical techniques, including regression analysis, step-wise linear regression, and correlation analysis, to evaluate our hypotheses. Our findings revealed that minimizing project-specific risks is important for reducing turnover intentions. Specifically, all project-specific risk factors, with the exception of control process risk, were associated with turnover intention through job satisfaction and influenced employee turnover. Additionally, while perceived job alternatives did not moderate the relationship between job satisfaction and turnover intention, they independently influenced turnover intention. Similarly, our analyses indicated that external and internal networking did not directly impact turnover intention. In summary, this study sheds light on critical factors influencing turnover intentions in Pakistan’s IT industry and underscores the importance of managing project-specific risks and perceived job alternatives to mitigate turnover.
@@ 11 Future Work
In the future, we plan to expand our research to explore project-specific risks across diverse industries, such as construction companies. By investigating the factors contributing to project-specific risks in various sectors, we aim to gain a comprehensive understanding of their influence on employee turnover intentions. Additionally, we intend to delve deeper into the determinants of employee turnover intentions, examining a wide range of influencing factors. This comprehensive approach will enable us to uncover detailed insights into the drivers of turnover intentions and their implications for organizational management. Moreover, we recognize the importance of considering cultural differences in our research. We aim to conduct cross-cultural analyses to explore how cultural norms and values shape employee turnover intentions. By collecting data from different cultural contexts, we can better understand the variations in turnover intentions across diverse populations.
@@
Furthermore, we plan to evaluate the outcomes of organizational initiatives and activities in mitigating employee turnover intentions. By assessing the impact of various organizational interventions, such as training programs, leadership initiatives, and employee engagement strategies, we can identify effective measures for reducing turnover and retaining skilled talent. Overall, our future research endeavors will contribute to the development of evidence-based strategies for minimizing employee turnover intentions and informing organizational outcomes, particularly in dynamic industries like IT where talent retention is relevant.
12 Ethical considerations
Participation was voluntary. No direct identifiers were collected. Responses were used only for research and reported in aggregate.
13 Appendix
13.1 Questionnaire
How much is your experience (years) in industry? *
  • < 5 years
  • 5–10 years
  • 11–15 years
  • 16–20 years
  • > 20 years
What is your salary range? *
  • < 25,000
  • 25,000–50,000
  • 51,000–75,000
  • 76,000–100,000
  • > 100,000
What is the company legal status? *
  • Public firm
  • Private firm
How many total employees your company has? *
Is your company a startup venture? *
  • Yes
  • No
Age *
  • Less than 25
  • 25–29
  • 30–39
  • 40–49
  • 50 or above
Please Rate the influence of project specific risks factor on your job satisfaction? *
a) Risks involved in Estimating software requirements
Unimportant Little Important Moderately Important Important Very Important
Delay in recruiting and resourcing $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Less or no experienced in similar project $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Estimation errors $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Inaccurate requirement analysis $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Miscommunication of requirements $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Conflicting and continuous requirements $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Language and regional differences with clients $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Lack of client ownership and owner stability $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Inaccurate cost measurement $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
b) Team Composition Risk *
Un important Little Important Moderately Important Important Very Important
Working with inexperienced team $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Team diversity $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Lack of availability of domain expert $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Lack of commitment from project team $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Lack of top management support $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Low morale of the team $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Please select the impact of branding employment and Peer perception regarding work on job satisfaction? *
Branding Employment means how people see the worth of job oppourtunity in the organization. Peer Perception Regarding Work means how your friends and other peers value you based upon your work
Strongly Disagree Disagree Neutral Agree Strongly Agree
External friends (like university friends and other friends of your field which are not working in your organization) finds your work interesting and innovative $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Internal peers sees tasks assigned to you interesting and innovative $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
You have a notable position in the community based upon the perception regarding your work $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Work experience of your organization is given more preference as compared to work experience of others $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Your company creates a sense of urgency and excitement about working at the company $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Your organization engages mind, heart and dreams $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Your organization provides a clear and compelling reason to work $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Your company is consistent with what employee believes about working at the organization. $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Your company evokes feeling of fun, prestige, challenge and rewards $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Please choose the importance of organization adaptation to changing technology and career opportunities on the job satisfaction *
a) Career opportunities
Strongly Disagree Disagree Neutral Agree Strongly Agree
Adequate work fulfillment $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Work useful enough $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Good use of your skills $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Chances to improve skills $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Promotional opportunities $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Lateral transfer opportunities $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Independence on the job $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Providing good quality on-the-job training $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Good chance of getting formal training $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Good quality formal training $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
b) Organization adaptation to changing technology
Strongly Disagree Disagree Neutral Agree Strongly Agree
Adaptation of cutting edge commonly used technology for company products and services $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Providing adequate training programs for commonly used technology used in company line of business $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Adaptation of latest principles, mechanisms and techniques for improving productivity and quality $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Advance planning for adaptation of futuristic technology as compared to other organizations $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Adequate financial resources required for quick change to new technology $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Please choose the importance and rate the influence of external social network, external and internal networking on perceived job alternatives
External Social Network:
Yes No
a) Are you a member of any professional association? * $\circ$ $\circ$
b) I regularly attend professional association meetings at the local level * $\circ$ $\circ$
c) I systematically read professional journals * $\circ$ $\circ$
d) Have you joined professional network site (e.g., LinkedIn)? * $\circ$ $\circ$
Yes No
e) Do you regularly maintain your profile on professional network site (e.g., LinkedIn)? * $\circ$ $\circ$
External networking *
Strongly Disagree Disagree Neutral Agree Strongly Agree
I spend a lot of time and effort at work networking outside my organization $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I am good at building relationships with influential people outside my organization $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I know a lot of important people and am well connected outside my organization $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I spend a lot of time developing connections with others outside my organization $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I am good at using my connections and network outside my organization to make things happen for my career $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I have developed a large network of colleagues and associates outside my organization that I can call on for support when I really need to get things done $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Internal networking *
Strongly Disagree Disagree Neutral Agree Strongly Agree
I spend a lot of time and effort networking with others within my organization $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I am good at building relationships with influential people within my organization $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I know a lot of important people and am well connected within my organization $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I spend a lot of time developing connections with others within my organization $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I am good at using my connections and network within my organization to make things happen at work $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
I have developed a large network of colleagues and associates within my organization whom I can call on for support when I really need to get things done $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Please choose how perceived job alternatives mediates or moderates the relationship between job satisfaction and employee turnover intentions.
Perceived Job Alternative *
Strongly Disagree Disagree Neutral Agree Strongly Agree
a) How easy would it be for you to find a job with another employer in this geographical area that is as good as the one you now have? $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
b) How easy would it be for you to find a job with another employer in this geographical area that is better than the one you now have? $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Please choose how perceived job alternatives mediates or moderates the relationship between job satisfaction and employee turnover intentions.
Perceived Job Alternative *
Strongly Disagree Disagree Neutral Agree Strongly Agree
a) How easy would it be for you to find a job with another employer in this geographical area that is as good as the one you now have? $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
b) How easy would it be for you to find a job with another employer in this geographical area that is better than the one you now have? $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
Job Satisfaction *
Strongly Disagree Disagree Neutral Agree Strongly Agree
In general, I like my job $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
In general, I am satisfied with my job $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
In general, I like working here $\circ$ $\circ$ $\circ$ $\circ$ $\circ$